### v1

# Synthesizing Sheet Music Problems for Evaluation and Reinforcement Learning

Abstract
Enhancing the ability of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) to interpret sheet music is a crucial step toward building AI musicians. However, current research lacks both evaluation benchmarks and training data for sheet music reasoning. To address this, we propose the idea of synthesizing sheet music problems grounded in music theory, which can serve both as evaluation benchmarks and as training data for reinforcement learning with verifiable rewards (RLVR). We introduce a data synthesis framework that generates verifiable sheet music questions in both textual and visual modalities, leading to the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench) and a complementary training set. Evaluation results on SSMR-Bench show the importance of models’ reasoning abilities in interpreting sheet music. At the same time, the poor performance of Gemini 2.5-Pro highlights the challenges that MLLMs still face in interpreting sheet music in a visual format. By leveraging synthetic data for RLVR, Qwen3-8B-Base and Qwen2.5-VL-Instruct achieve improvements on the SSMR-Bench. Besides, the trained Qwen3-8B-Base surpasses GPT-4 in overall performance on MusicTheoryBench and achieves reasoning performance comparable to GPT-4 with the strategies of Role play and Chain-of-Thought. Notably, its performance on math problems also improves relative to the original Qwen3-8B-Base. Furthermore, our results show that the enhanced reasoning ability can also facilitate music composition. In conclusion, we are the first to propose the idea of synthesizing sheet music problems based on music theory rules, and demonstrate its effectiveness not only in advancing model reasoning for sheet music understanding but also in unlocking new possibilities for AI-assisted music creation. Our code and resources are available at GitHub.
1 Introduction
"The sheet music is the language of musicians."
Recent advancements in Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have inspired researchers to explore the potential of developing AI musicians (Qu et al., 2025; Bradshaw & Colton, 2025; Wang et al., 2024). Given that sheet music is the universal language of musicians, the ability to read and interpret it is an essential step for AI musicians  (Yuan et al., 2024; Wang et al., 2025). This requires a grasp of the core concepts of music theory, which provide the structural and expressive framework necessary for musical interpretation. Enhancing LLMs and MLLMs to interpret sheet music involves not just recognizing musical symbols but also modeling the cognitive processes involved in understanding music.
However, research on sheet music reasoning remains limited, with a lack of standardized evaluation benchmarks and training datasets. For example, ChatMusician (Yuan et al., 2024) introduces MusicTheoryBench, a benchmark consisting of 367 human-designed questions, some of which specifically target sheet music understanding in ABC notation, a textual format of sheet music. MMMU (Yue et al., 2024) provides a small set of sheet music understanding questions in image format, though only 30 of them are publicly released with answers. In contrast, MusiXQA (Chen et al., 2025a) contains a larger set of image-based sheet music data, but it mainly targets the recognition of individual elements rather than reasoning about their connections within music theory.
Inspired by recent human-programmatic approaches to data synthesis (Liu et al., 2025; Chen et al., 2025b; Li et al., 2025), we propose to leverage the rules of music theory to programmatically generate an endless supply of verifiable reasoning problems about sheet music. Such a resource is invaluable for creating robust evaluation benchmarks and for training models with verifiable rewards (Shao et al., 2024; Yu et al., 2025) to improve their reasoning abilities. The practical implementation of this concept is our novel data synthesis framework. This system is designed to generate questions with staff notation presented in both textual and visual formats and includes nine Question Template Classes, categorized into Rhythm, Chord, Interval, and Scale. We use this framework to build the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench). The benchmark includes an evaluation set with 1,600 textual and 1,600 visual QA pairs, and a larger training set with 8,000 pairs for each modality.
We evaluate SSMR-Bench using a range of state-of-the-art LLMs and MLLMs. For the textual QA task, which represents sheet music in ABC notation, Qwen3-32B demonstrates performance on par with Deepseek-V3, while Deepseek-R1 exceeds Deepseek-V3 by 18.69% overall. This highlights the model’s enhanced reasoning capability in understanding sheet music. For the visual QA task, Gemini2.5-Pro achieves the best performance with an average score of 56.62%, surpassing Qwen2.5-VL-72B-Instruct, InternVL3-78B, and other models. However, this performance is significantly lower than when the same questions are presented in textual format, highlighting the challenges of interpreting sheet music from images.
We train Qwen3-8B-Base and Qwen2.5-VL-7B-Instruct on our synthetic dataset using Group Relative Policy Optimization (GRPO,  Shao et al., 2024). Both models achieve substantial improvements. Specifically, the trained Qwen3-8B-Base achieved a 13.06% improvement on our SSMR-Bench over Qwen3-8B, reaching performance comparable to GPT-4.1-mini. Similarly, the trained Qwen2.5-VL-7B-Instruct improves its ability to recognize and analyze sheet music from images, allowing it to outperform Gemini2.5-Pro. To verify the effectiveness of our synthetic data, we further evaluate the trained Qwen3-8B-Base on MusicTheoryBench (Yuan et al., 2024). The results show that the model outperforms GPT-4 in overall performance on MusicTheoryBench and achieves reasoning performance comparable to GPT-4 with Role Play and Chain-of-Thought. Besides, training Qwen3-8B-Base on our synthetic data also boosts its performance on math problems, surpassing the original model’s capabilities. Furthermore, the improved reasoning ability in sheet music understanding can also facilitate music composition.
In conclusion, we summarize our key contributions as follows:
  • We are the first to propose the idea of leveraging music theory rules to programmatically synthesize verifiable sheet music problems, using them as both an evaluation benchmark and a training set for Reinforcement Learning with Verifiable Rewards (RLVR).
  • We develop a data synthesis framework based on this idea, which is capable of generating verifiable sheet music questions with staff notation in both textual and visual modalities, and we provide SSMR-Bench along with a corresponding training set in both modalities.
  • We conduct comprehensive experiments to demonstrate the potential of automatically synthesizing verifiable sheet music problems based on music theory rules. Training on this synthetic data enhances models’ reasoning abilities in sheet music, and this improved ability also shows promise in facilitating music composition.
2 Related Work
2.1 Sheet Music Understanding
The application of artificial intelligence to sheet music analysis has consistently been a focal point for the research community (Fornés et al., 2012; Shatri & Fazekas, 2024). Optical Music Recognition (OMR) employs AI-driven tools to interpret music notation, necessitating both specialized expertise and domain knowledge (Ma et al., 2024; Tuggener et al., 2018; Calvo-Zaragoza & Rizo, 2018). Some research in this area focuses on the recognition of handwritten music scores, introducing widely used benchmarks and datasets (Torras et al., 2024; Mayer et al., 2024), while other studies propose novel methodologies to address the underlying challenges (Rebelo et al., 2012; Rıos-Vila et al., 2024; R´ıos-Vila et al., 2024; Li et al., 2023).
While significant progress has been made, some researchers are expanding the study to Sheet Music Question Answering (QA), moving beyond traditional music recognition and image modalities. MMMU (Yue et al., 2024) provides music QA, highlighting the challenges of sheet music QA based on images. Meanwhile, there is growing interest in using LLMs for symbolic music (Wu et al., 2024; Wang et al., 2025; Qu et al., 2025), where sheet music is represented in textual formats, such as the widely used ABC notation.  Yuan et al. (2024) introduces a MusicTheoryBench that includes knowledge and reasoning categories to evaluate the music understanding abilities of LLMs. However, the manually crafted nature of the two benchmarks hinders data scalability, which is the problem we aim to address in this work.
2.2 Reinforcement Learning Through Synthetic Data
Reinforcement Learning (RL) has proven to be a critical factor in enhancing the reasoning capabilities of models (DeepSeek-AI et al., 2025; OpenAI et al., 2024). Following the success of DeepSeek-R1, recent work has shifted towards Reinforcement Learning with Verifiable Rewards (RLVR), which utilizes objectively verifiable rewards derived from objective questions and their corresponding answers (Shao et al., 2024; Yu et al., 2025; Yan et al., 2025). For RLVR, how to obtain the verifiable data is a crucial question. Recently, some research has attempted to address this challenge by synthesizing data (Li et al., 2025). Enigmata (Chen et al., 2025b) provides fully synthesizable data across 36 puzzle problems, while SynLogic (Liu et al., 2025) proposes a data synthesis framework that generates diverse logical reasoning data, covering 35 distinct tasks. Compared to these domains, music is a distinct field, yet it still follows common rules that can be applied to design verifiable questions (Perricone, 2018; Mulholland & Hojnacki, 2013; Terefenko, 2017), which is the core focus of this work.
3 Synthesize Music Reasoning Questions
3.1 Data Synthesize Framework
Figure 1: Overview of synthesizing verifiable sheet music questions. Experts define rule-based Question Template Classes that generate questions and answers from ABC notation, where the music sheet can be converted into an image, enabling the construction of VQA-style questions. The process is fully programmatic, without LLMs.
Music is traditionally written on a staff, a set of five lines where visual symbols represent the core elements of music theory. The vertical position of a notehead indicates its pitch, while its shape (e.g., hollow or filled, with a stem or flag) defines its rhythm, or duration. These visual cues are organized into measures, separated by barlines, according to a time signature that dictates the underlying rhythmic pulse. While staff notation is intuitive for human musicians, its graphical nature is cumbersome for programmatic analysis. To bridge this gap, we utilize ABC notation, a text-based format that encodes the same musical information in a machine-readable way. In ABC, pitches are represented by letters (C, D, E), rhythmic values are specified with numerical modifiers, and measures are delineated by vertical bars (|). This format allows us to translate the abstract rules of music theory into parsable strings. Examples of ABC notation are provided in Appendix A.
Our framework is designed to generate questions that test a model’s understanding of these fundamental principles, from rhythmic calculation and chord identification to the analysis of intervals and scales. We outline the process of synthesizing sheet music questions based on music theory rules in Figure 1. Sheet music is first obtained in ABC notation. Then, the experts design specialized Question Template Classes derived from music theory rules, each capable of generating both questions and corresponding answers from the ABC representation. Applying these templates produces textual QA pairs. Subsequently, the sheet music in each question can be converted into images using abcm2ps111abcm2ps converts ABC notation to PostScript sheet music. https://github.com/lewdlime/abcm2ps and ImageMagick [^2] Tool for converting images, e.g., PS or SVG to JPG. https://github.com/ImageMagick/ImageMagick, enabling the construction of VQA questions. The entire pipeline is fully programmatic, without reliance on LLMs, ensuring controllability and interpretability.
3.2 Data Construction
The scalable synthesis framework we designed currently provides 9 types of questions. All questions are categorized into four categories: Rhythm, Chord, Interval, and Scale, which are the fundamental elements of sheet music. The cases for each question type are provided in Appendix A.2.
Figure 2: The composition of the SSMR-Bench test set consists of four categories: Rhythm, Chord, Interval, and Scale. Each category contributes equally, accounting for 25% of the total.
  • Rhythm Questions challenge models to understand the rhythm of sheet music. These questions require models to calculate note values and analyze the music from a holistic perspective.
  • Chord Questions assess models’ reasoning ability to identify and infer chord structures from given notes, testing their understanding of harmonic relationships and tonal context.
  • Interval Questions evaluate the model’s ability to recognize and compute the distance between two notes in terms of pitch, requiring a nuanced understanding of musical intervals.
  • Scale Questions test the model’s proficiency in identifying scales and key signatures, analyzing the relationships between notes and their positions within different scale types.
To solve these questions, LLMs and MLLMs need basic music knowledge, sheet music reading skills, and reasoning ability.
Through our data synthesis framework, we introduce the Synthesize Sheet Music Reasoning Benchmark (SSMR-Bench), comprising 1,600 textual and 1,600 visual questions. The sheet music used to generate questions is sampled from MelodyHub (Wu et al., 2024), a large-scale dataset in ABC notation. The distribution of question types and categories within the benchmark is illustrated in Figure 2. Furthermore, we offer a comprehensive training set, consisting of 8,000 textual and 8,000 visual questions. A comparative analysis of our framework with prior music sheet-related benchmarks is presented in Table 1. Our framework produces both visual and textual modality questions for music sheet reasoning and constructs a verifiable training set without human curation, thereby enabling scalable data generation.
Table 1: Comparison of music-related QA datasets across different aspects. Checkmarks (✓) denote presence, crosses (✗) absence, and triangles ($\triangle$) partial coverage.
Dataset Sheet Music QA Synthetic Modality Trainable
Textual Visual
MMMU (Yue et al., 2024)
MusiXQA (Chen et al., 2025a)
ZIQI-Eval (Li et al., 2024)
MusicTheoryBench (Yuan et al., 2024) $\triangle$
Ours
4 Experiment
4.1 Experimental Setup
Evaluation.
We select the frontier LLMs and MLLMs to evaluate their performance on our SSMR-bench. For testing, the temperature is set to 0.7, with a length limit of 8192. We shuffle the multiple-choice options to avoid contamination. We utilize Math-Verify to parse the answers from the model responses and vLLM as the backend for inference serving.
RL Practice.
We employ Qwen3-8B-Base for reinforcement learning in the textual modality and Qwen2.5-VL-7B-Instruct for the visual modality. In addition, we include a variant of Qwen3-8B-Base trained on 8,000 math problems sampled from OpenR1 (Hugging Face, 2025) to compare against the effectiveness of synthetic music reasoning data. Training is performed using the GRPO algorithm with a batch size of 128, a rollout number of 8, and an update batch size of 64. The KL loss and entropy loss are disabled by setting their coefficients to zero. The reward function relies solely on accuracy, without incorporating format-related signals. The training template is provided in Appendix B. The total number of training steps is 300.
4.2 Results on Synthetic Music Reasoning Benchmark
Table 2 presents the performance of advanced models on SSMR-bench, covering both LLMs and MLLMs in the textual and visual modalities. The performance metrics reveal notable variations between different models, highlighting their strengths and limitations in each modality.
Results on Textual Modality.
The results for textual QA show that the LLMs, such as Deepseek-R1, outperform others significantly across all categories. Specifically, Deepseek-R1 achieves the highest overall accuracy of 93.63%, with exceptional performance in Scale (99.75%). In contrast, models like Qwen2.5-32B-Instruct and Llama3.1-70B-Instruct exhibit lower accuracies, with scores of 51.25% and 43.94%, respectively. Llama3.1-8B-Instruct shows the lowest performance with an overall score of 20.94%.
Results on Visual Modality.
The performance of MLLMs on sheet music reasoning, with staff notation presented in image format, is substantially lower than that of LLMs on textual reasoning. The best-performing MLLM is Gemini 2.5 Pro, which is recognized as one of the most advanced models available. It achieves an overall accuracy of 55.44%, with particularly strong performance on Scale (84.75%). Other models, such as Qwen2.5-VL-72B-Instruct, InternVL3-78B, and InternVL2.5-78B, show accuracies of 45.56%, 45.75%, and 40.44%, respectively.
Comparison Between Modality.
The lower performance of MLLMs in the visual modality can be attributed to the unique challenges posed by music element recognition. Unlike textual data, which is relatively machine-friendly, musical staff images contain intricate visual patterns. To interpret these images, MLLMs must first recognize diverse musical elements, such as notes, clefs, time signatures, and other notations, before reasoning about their relationships to answer specific questions. This dual process presents two major challenges: (1) accurately identifying musical elements in the image, and (2) reasoning about these elements to generate appropriate answers.
Effectiveness of Reasoning.
The incorporation of reasoning capabilities significantly enhances model performance on music reasoning tasks. Notably, models like Qwen3-32B and QwQ-32B, with only 32 billion parameters, achieve results comparable to Deepseek-V3, which has 671 billion parameters. Moreover, Deepseek-R1, a reasoning model derived from Deepseek-V3, outperforms Deepseek-V3 by 18.69% in overall performance. This underscores the critical role of reasoning in boosting model accuracy and highlights that even models with fewer parameters can achieve competitive performance when equipped with advanced reasoning capabilities.
Table 2: Evaluation results of frontier models for textual and visual QA on the SSMR-bench. The table presents the performance of various models across different sheet music reasoning tasks, including Rhythm, Chord, Interval, and Scale, with the overall performance summarized for both textual and visual QA.
Models Synthesize Sheet Music Reasoning Benchmark
Rhythm Chord Interval Scale Overall
$(400)$ $(400)$ $(400)$ $(400)$ $(1600)$
Random choice $25.00$ $25.00$ $25.00$ $25.00$ $25.00$
Textual QA
Deepseek-R1 $93.00$ $90.50$ $91.25$ $99.75$ $93.63$
Qwen3-32B $80.00$ $75.00$ $62.25$ $89.50$ $76.69$
Deepseek-V3-250324 $78.75$ $71.75$ $58.75$ $90.50$ $74.94$
QwQ-32B $69.75$ $72.50$ $59.50$ $87.75$ $72.38$
GPT-4.1-mini $77.00$ $69.75$ $56.25$ $83.75$ $71.69$
Qwen3-8B $54.00$ $60.50$ $50.75$ $66.25$ $57.88$
Qwen2.5-32B-Instruct $54.50$ $47.00$ $53.75$ $49.75$ $51.25$
GPT-4o-mini $48.00$ $54.50$ $48.50$ $40.25$ $47.81$
Llama3.1-70B-Instruct $39.75$ $41.00$ $47.00$ $48.00$ $43.94$
Llama3.1-8B-Instruct $23.50$ $19.00$ $18.00$ $23.25$ $20.94$
Visual QA
Gemini 2.5 Pro $44.75$ $61.25$ $32.25$ $83.50$ $55.44$
InternVL3-78B $34.25$ $44.75$ $29.00$ $75.00$ $45.75$
Qwen2.5-VL-72B-Instruct $39.75$ $41.75$ $27.50$ $73.25$ $45.56$
InternVL2.5-78B $29.00$ $31.00$ $32.25$ $69.50$ $40.44$
4.3 RLVR on Synthetic Music Reasoning Data
Compared to previous approaches, our data synthesis framework provides a simple yet effective method for scaling sheet music reasoning data with verifiable answers, enabling RLVR on LLMs and MLLMs. We train two distinct models on our synthetic dataset of 8,000 examples. The text-based Qwen3-8B-Base is trained on textual data, while the vision-language model, Qwen2.5-VL-7B-Instruct, is trained on visual data. Both models are optimized using the GRPO algorithm. Table 3 presents the performance of these models on our SSMR-Bench.
In the Textual Modality, the Qwen3-8B-Base model initially achieved relatively low scores across all metrics, with an average score of 23.18%, prior to the application of GRPO. However, after incorporating GRPO on the synthetic training set, performance shows significant improvements, with average scores increasing by 47.76% points, resulting in an impressive final score of 70.94%. This also represents a 13.06% improvement over the baseline Qwen3-8B score of 57.88%. In the Visual Modality, the Qwen2.5-VL-7B-Instruct model initially recorded moderate results with an average score of 41.81%, prior to applying RLVR. Following RL training, the model’s performance improved by an average of 28.75% points, reaching a final score of 70.56%.
Figure 3 illustrates the average rewards and response length changes throughout the model training process. As shown in the left panel, both models demonstrate a steady increase in rewards, indicating effective learning from the reward signals. Notably, the Qwen3-8B-Base model, trained with textual modality data, consistently outperforms the Qwen2.5-VL-7B-Instruct model, which uses visual data. This difference can be attributed not only to the models’ inherent capabilities but also to the higher difficulty associated with VQA. The right panel illustrates the trends in response length. The Qwen3-8B-Base model starts with longer responses, which quickly drop to around 350 tokens, then gradually and steadily increase to 500 tokens. In contrast, the Qwen2.5-VL-7B-Instruct model begins with shorter responses, decreases to 150 tokens, and then gradually increases, exhibiting greater variability in the later stages of training, eventually reaching 800 tokens. This divergence suggests that the optimization process drives the models toward different response lengths, influenced by their inherent abilities and respective modalities. Although both models exhibit some divergence, the overall trend during training remains consistent.
Table 3: Performance of Qwen3-8B-Base and Qwen2.5-VL-7B-Instruct on the SSMR-bench after applying GRPO on synthetic data, with results compared to the Qwen3-8B and Qwen2.5-VL-7B-Instruct, respectively. The table also highlights the improvements.
Models Synthesize Sheet Music Reasoning Benchmark
Rhythm Chord Interval Scale Overall
Textual QA
Qwen3-8B-Base $23.25$ $32.25$ $18.25$ $19.00$ $23.18$
    + GRPO 73.7573.75(+50.50) 59.0059.00(+26.75) 62.0062.00(+43.75) 89.0089.00(+70.00) 70.9470.94(+47.76)
Qwen3-8B $54.00$ $60.50$ $50.75$ $66.25$ $57.88$
Visual QA
Qwen2.5-VL-7B-Instruct $44.00$ $40.75$ $35.25$ $47.25$ $41.81$
    + GRPO 66.5066.50(+22.50) 68.7568.75(+28.00) 55.2555.25(+20.00) 91.7591.75(+44.50) 70.5670.56(+28.75)
Figure 3: Training dynamics of GRPO on textual and visual modality. (Left) The reward curves for both Qwen3-8B-Base + Textual GRPO and Qwen2.5-VL-7B-Instruct + Visual GRPO show a steady increase, indicating that both models are effectively learning from the reward signal. (Right) The average response length of both models initially decreases and then increases. While the Qwen3-8B-Base model shows a steady increase, the response length of Qwen2.5-VL-7B-Instruct experiences notable fluctuations in the later stages of training.
4.4 Results Analysis by Question Type
Figure 4: Performance comparison of various selected LLMs and MLLMs on different music-related question types. The upper plot shows the performance of LLMs evaluated on the textual modality, while the lower plot displays the performance of MLLMs in the visual modality. The performance is assessed across nine distinct question categories, such as Bar Placement, Chord Identification (Chord ID), and Scale Identification (Scale ID). Detailed descriptions of these question types can be found in Appendix A.
To gain deeper insights into the model’s performance, we select various LLMs and MLLMs and present their performance categorized by question type. We present the results in Figure 4.
The upper plot of Figure 4 illustrates the performance of LLMs in the textual modality. After reinforcement learning on 8,000 synthetic music reasoning samples through GRPO, Qwen3-8B-Base outperforms Qwen3-8B in 2/3 of the question types, despite the latter undergoing extensive data training. Additionally, Qwen3-8B-Base+GRPO achieves comparable performance to Deepseek-V3 and Qwen3-32B in aspects such as "Time Sig", "Interval No", and "Note Comp". However, for "Chord ID" and "Scale Sel," the model shows limited improvement, which can be attributed to the scarcity of these data types in the training set shown in Appendix A.
The lower plot of Figure 4 illustrates the performance of various selected MLLMs in the visual modality. The results highlight the superior performance of the Qwen2.5-VL-7B-Instruct+GRPO model, which consistently achieves the highest accuracy across all question types. This improvement stems from reinforcement learning, which enhances the ability of Qwen2.5-VL-7B-Instruct to accurately recognize elements of sheet music in image format and perform reasoning based on these visual cues.
5 Beyond Synthetic Data and Question Answering Tasks
5.1 RL on Synthetic Data Benefits Real-World Scenarios
To assess the effectiveness of our synthetic data, we conduct experiments on MusicTheoryBench (Yuan et al., 2024). Specifically, we evaluate the performance of Qwen3-8B-Base after GRPO on the synthetic sheet music problems and compare it against three baselines: the unmodified Qwen3-8B-Base, the officially released reasoning model Qwen3-8B, and a GRPO-trained variant using MATH data. Additionally, we report results for GPT-4, GPT-4 with Role Play, and GPT-4 with Chain-of-Thought (CoT) reasoning (Yuan et al., 2024) to provide further context. The results are summarized in Table 4. Besides, we also evaluate the Qwen3-8B-Base model and its variants on a series of math benchmarks, and present results in Table 5.
As shown in Table 4, Qwen3-8B-Base demonstrates substantial performance gains after GRPO training on the synthetic data. Its original performance was 37.92% in knowledge and 19.39% in reasoning, averaging 28.65%. After GRPO, the model improves by 25.28% in knowledge and 17.34% in reasoning, corresponding to an average gain of 21.32%. Compared to Qwen3-8B, the improvement is mainly observed in reasoning (up by 7.14%) with a marginal increase in knowledge (0.37%). In contrast, the Qwen3-8B-Base variant trained with GRPO on MATH data exhibits much smaller gains, achieving only a 33.81% average score, which is 12.4% lower than Qwen3-8B. Besides, Qwen3-8B-Base reinforced with our synthetic data outperforms GPT-4 in overall performance and achieves reasoning capabilities comparable to GPT-4 with Role Play and GPT-4 with CoT.
In the mathematical domain, training the Qwen3-8B-Base model on our synthetic dataset yields substantial improvements. After training, the model demonstrates enhanced accuracy across multiple benchmarks. For example, its performance on the AIME24 dataset increases from 8.75% to 13.75%, while on the AMC dataset, accuracy rises from 39.16% to 52.11%. The model also demonstrates significant improvements on Minerva, increasing from 25.37% to 34.19%, and on Olympiad Bench, rising from 30.22% to 40.44%. Moreover, when evaluated on the MATH-500 dataset, the Qwen3-8B-Base+Music RL model achieves a remarkable accuracy of 79.40%. Nevertheless, this performance remains significantly lower than that of the version trained on math-specific data, highlighting the critical importance of domain-specific datasets.
This confirms the effectiveness of synthesizing verifiable questions specifically related to sheet music in enhancing the model’s understanding and reasoning abilities within this domain, while yielding only moderate gains in Math problem-solving. In contrast, applying GRPO to MATH data yields substantially smaller gains in sheet music reasoning while producing significant improvements in mathematical tasks, underscoring the critical importance of domain relevance in synthetic data generation. Overall, these findings suggest that automatically generating verifiable sheet music problems provides a practical and effective approach for strengthening reasoning capabilities in sheet music comprehension.
Table 4: Evaluation results on the MusicTheoryBench. Training the Qwen3-8B-Base model with GRPO and our synthetic data leads to substantial gains in both knowledge and reasoning. Notably, our enhanced model also outperforms the Qwen3-8B, especially in the reasoning category.
Models MusicTheoryBench
Knowledge Reasoning Avg.
GPT4-0-shot $58.20$ $25.60$ $41.90$
GPT4-CoT $68.40$ $36.70$ $52.55$
GPT4-RolePlay $68.30$ $36.60$ $52.45$
Qwen3-8B $62.83$ $29.59$ $46.21$
Qwen3-8B-Base $37.92$ $19.39$ $28.65$
   + Math GRPO 47.2147.21(+9.29) 20.0420.04(+0.65) 33.8133.81(+5.16)
   + Music GRPO 63.2063.20(+25.28) 36.7336.73(+17.34) 49.9749.97(+21.32)
Table 5: Performance of Qwen3-8B series models on various math benchmarks, with results after applying Music RL and Math RL compared to baselines.
Models Math Benchmarks
AIME24 AMC MATH-500 Minerva Olympiad Bench
Qwen3-8B $50.41$ $71.08$ $91.60$ $45.96$ $60.89$
Qwen3-8B-Base $8.75$ $39.16$ $66.00$ $25.37$ $30.22$
+ Math GRPO 25.8325.83(+17.08) 62.0462.04(+22.88) 85.4085.40(+19.40) 40.4440.44(+15.07) 50.5250.52(+20.30)
+ Music GRPO 13.7513.75(+5.00) 52.1152.11(+12.95) 79.4079.40(+13.40) 34.1934.19(+8.82) 40.4440.44(+10.22)
5.2 From Question Answering to Sheet Music Generation
To investigate whether improved sheet music reasoning can facilitate music composition, we extend our evaluation from question answering to sheet music generation. Specifically, we prompt models to generate a four-measure continuation conditioned on an initial four-measure excerpt of sheet music in ABC notation. We evaluate these continuations on rhythmic consistency, a dimension that prior studies have identified as particularly challenging for sheet music generation (Suzuki, 2021).
We introduce two primary evaluation metrics:
  • Rhythmic Consistency (RC): The average of a binary score assigned to each generated continuation. A sample is scored 1 if all four measures are rhythmically correct and 0 otherwise.
  • Novel Rhythmic Coherence (NRC): While melody repetition is common in music, we also aim to assess a model’s likelihood of producing rhythmically correct sheet music when generating novel melodies. The NRC metric is designed for this purpose. It is calculated as the RC score, but only for generated continuations that are not direct copies of the provided prompt.
Figure 5: Comparison of rhythmic consistency. The figure shows the Rhythmic Consistency (RC) and Novel Rhythmic Coherence (NRC) for Qwen3-8B-Base and its GRPO variant, trained on our synthetic data.
We randomly sample 30 pieces of sheet music from SSMR-bench and evaluate each model’s continuation through human judgment. The results in Table 5 show that training on synthetic sheet music problems also leads to improvements in music composition. The base Qwen3-8B model achieves an RC score of 66.67%. After reinforcement learning on our synthetic dataset, this score dramatically increases to 99.33%, surpassing the 70.00% from the original Qwen3-8B. The improvement in generating novel content is also evident: our fine-tuned model’s NRC score of 88.89% is a significant increase over the base model’s 50.00% and the original Qwen3-8B’s 68.97%. Although we focus only on the rhythm dimension, which is just one aspect of music, the observed improvements also highlight the potential for enhancing reasoning abilities to facilitate music composition.
6 Discussion
Possible Challenges Facing AI Musicians.
The findings presented in Section 4.2 highlight the varying levels of difficulty associated with different modalities of sheet music. In particular, the same question, when presented in visual sheet music format, is significantly more challenging for LLMs than when conveyed in a textual format. Beyond modality, sheet music can be represented in a variety of formats, including MIDI, MusicXML, and others. Furthermore, sheet music is not limited to traditional staff notation; alternative formats, such as Guitar Tablature and numbered musical notation, should also be considered, which together present substantial challenges for the development of an ideal AI musician.
Limitations and Future Work.
In this work, we focus on synthesizing verifiable sheet music reasoning problems based on widely accepted rules in music theory. We implement this approach through a data generation framework and evaluate its effectiveness. However, the questions we construct are relatively simple, and can be reliably solved by Deepseek-R1. At the same time, they highlight the challenge for MLLMs to fully understand music sheets, as evidenced by the low performance of Gemini2.5-Pro. Taken together, these observations motivate future work on designing more complex, rule-based, verifiable music reasoning questions and on developing improved methods to enhance models’ reasoning abilities of sheet music, as well as transferring these abilities to music composition.
7 Conclusion
In this work, we introduce the concept of synthesizing sheet music problems based on the rules of music theory, enabling the construction of verifiable questions for evaluation and training. We implement this idea by proposing a framework that automatically generates sheet music questions with staff notation in both textual and visual modalities. Based on this framework, we introduce SSMR-Bench and a corresponding training set in both modalities. Experimental results demonstrate that the synthesized data effectively enhances the reasoning capabilities of LLMs and MLLMs in sheet music tasks, underscoring its potential as a valuable resource. Moreover, the results also highlight the promise of this enhanced ability to facilitate music composition. Our work demonstrates the potential of programmatically generating verifiable sheet music problems to evaluate and improve the abilities of LLMs and MLLMs in sheet music, while also highlighting promising avenues for future research toward AI musicianship.
Appendix A Dataset Details
Table 6: Statistics of the question dataset, detailing the counts for each question class and its designated abbreviation. Questions are grouped into four primary music theory domains for both the Textual and Visual QA components. The dataset is balanced, containing 400 questions per domain.
Domain Question Class Abbreviation Counts Total
Textual QA
Scale ScaleIdentificationFromAbcQuestion Scale ID 352 400
ScaleSelectionQuestion Scale Sel 48
Rhythm TimeSignatureQuestion Time Sig 217 400
BarLinePlacementQuestion Bar Placement 183
Interval IntervalNumberQuestion Interval No 199 400
NoteCompletionByInterval Note Comp 201
Chord ChordsCompletionQuestion Chord Comp 156 400
ChordKeyRootIdentificationQuestion Chord Root ID 200
ChordIdentificationQuestion Chord ID 44
Visual QA
Scale ScaleIdentificationFromAbcQuestion Scale ID 352 400
ScaleSelectionQuestion Scale Sel 48
Rhythm TimeSignatureQuestion Time Sig 217 400
BarLinePlacementQuestion Bar Placement 183
Interval IntervalNumberQuestion Interval No 199 400
NoteCompletionByInterval Note Comp 201
Chord ChordsCompletionQuestion Chord Comp 156 400
ChordKeyRootIdentificationQuestion Chord Root ID 200
ChordIdentificationQuestion Chord ID 44
We provide a detailed description of the sheet music QA dataset used in our experiments in Table 6. The dataset is designed to evaluate a model’s reasoning ability in sheet music, across two distinct modalities: Textual QA and Visual QA.
The dataset is structured around four core domains of music theory: Scale, Rhythm, Interval, and Chord. It contains a total of 3,200 question-answer pairs, with a balanced distribution of 1,600 questions for each modality. Furthermore, each of the four domains is balanced with 400 questions per modality, ensuring that no single topic is over-represented.
A.1 Dataset Distribution
The precise distribution of questions across all classes and domains is provided in Table 6. The dataset was intentionally balanced at the domain level to ensure robust evaluation across different areas of music theory.
A.2 Question Class Descriptions
Below we describe the objective of each question class within the four domains. For Textual QA, the musical context is provided in ABC notation. For Visual QA, the context is a rendered image of standard music notation.
A.2.1 Domain: Scale
ScaleIdentificationFromAbcQuestion(Scale ID): Given a sequence of notes representing a scale, the model must identify the scale’s name.
[…]
ScaleSelectionQuestion(Scale Sel): Given a scale name and a note, the model must determine if the note belongs to that scale.
[…]
A.2.2 Domain: Rhythm
This domain focuses on rhythm, time signatures, and measure structure.
TimeSignatureQuestion (Time Sig): Given a complete measure of music, the model must deduce its time signature.
[…]
BarLinePlacementQuestion (Bar Placement): Given a time signature and an unbarred sequence of notes, the model must identify the correct position for the bar line.
[…]
A.2.3 Domain: Interval
This domain tests the ability to identify and construct intervals between notes.
IntervalNumberQuestion (Interval No): Given two notes, the model must identify the interval between them.
[…]
NoteCompletionByInterval (Note Comp): Given a starting note and an interval, the model must identify the resulting note.
[…]
A.2.4 Domain: Chord
This domain assesses knowledge of chord structure, identification, and composition.
ChordsCompletionQuestion (Chord Comp): Given two notes of a standard triad, the model must provide the missing third note.
[…]
ChordKeyRootIdentificationQuestion (Chord Root ID): Given a chord, the model must identify its root note.
[…]
ChordIdentificationQuestion (Chord ID): Given a set of notes, the model must identify the full name of the chord.
[…]
A.3 Data Format
Each sample in the dataset is a JSON object containing the unique identifier, modality, domain, question class, context, question, and the ground-truth answer. An example of a data point for the Textual QA modality is shown below.
[…]
For the Visual QA modality, the "context" field would contain a path to the corresponding image file (e.g., "image/visual-chords-0044.png").
Appendix B Training Template
All trained models employ an identical system prompt during both training and inference.
[…]